OpenStudy (anonymous):

The area of a rectangle is 72 square inches. If the length is 40 inches greater than four times the width. What are the dimensions of the rectangle. I need one variable in the equation!

OpenStudy (anonymous):

I need help with this problem please someone help me.

OpenStudy (campbell_st):

let the width be x the length is 4x + 40 the area x(4x + 40) = 72 expand and simplify 4x^2 + 40x - 72 = 0 4(x^2 + 10x - 18) = 0 use the general quadratic formula to solve... and only consider the positive value for x as distance can only be positive

OpenStudy (anonymous):

Solve the following for L and w: { L w = 72, L = 40 + 4 w }\[\left\{L=4 \left(5+\sqrt{43}\right),w=-5+\sqrt{43}\right\} \]